OpenStudy (anonymous):

Determine the sum for the positive two-digit integers that are not divisible by 5. A. 3960 B. 1640 C. 2830 D. 4550

OpenStudy (anonymous):

Determine the sum for the positive two-digit integers that are not divisible by 5. 11, 12, 13, 14, 16, 17, 18, 19, 21 ... 99 lets do this S(90) = (90/2)(11 + 99) S(90) = 4950 but were not done yet 10, 15, 20, 25, 30, 35, ... 95 S(9) = (9/2)(10 + 90) + (9/2)(15 + 95) S(9) = (18/2)(100 + 110) S(9) = 1890 now we just subtract 1890 from 4950 ANS : 3060

OpenStudy (anonymous):

3060 is not one of the choices

OpenStudy (anonymous):

I think its A

OpenStudy (raden):

it means : (10+11+12+13+...+99) - (10+15+20+25+...+95) = ...

OpenStudy (anonymous):

Okay